3x+7y=-4
-7x+y=44

Answer by mananth(16946)

3x+7y=-4............1
-7x+y=44.............2
Eliminate y
multiply (1)by 1
Multiply (2) by -7
3x+7y=-4
49x-7y=-308
Add the two equations
52x=-312
/52
x=-6
plug value of x in (1)
3x+7y=-4
-18+7y= -4
7y=-4+18
7y=14
y=2